### v2

@@ 4.1 Extensive & Intensive Margins
In Table 4, we present regression results that examine the extensive and intensive margins of labor supply for same-sex partnerships. The coefficient $\beta_{1}$ captures the effect of anti-discrimination laws on the labor supply outcomes for men in male same-sex partnerships and for women in female same-sex partnerships.
@@
Panel A focuses on men in male same‐sex partnerships. Columns (1) and (2) show that anti-discrimination laws reduce both the labor force participation and employment gaps by approximately 3.8 and 3.7 percentage points, which translates into reductions of about 78% and 77% respectively. These coefficients are statistically significant at the 1% level. Columns (3) through (6) address the intensive margin, indicating modest increases in weekly hours and weeks worked among all men. However, when the sample is restricted to employed men (Columns (5) and (6)), these effects become statistically insignificant. This pattern suggests that the primary channel through which these laws operate is by increasing the rate of employment.
Panel B reports analogous estimates for women in female same‐sex partnerships. In contrast, the extensive margin results for women (Columns (1) and (2)) yield statistically insignificant coefficients of $-0.37$ and $-0.50$ percentage points. Women in same‐sex partnerships exhibit higher labor force participation rates than women in different‐sex partnerships. This higher baseline may partially explain the lack of significant improvement on the extensive margin, as there is less room for gains. On the intensive margin, Column (3) shows a modest increase of 0.95 weekly hours (significant at the 10% level) for all women, while Column (4) finds a small, statistically insignificant reduction in weeks worked. When focusing on employed women (Columns (5) and (6)), we observe a statistically significant increase of 0.70 weekly hours (at the 5% level) alongside a significant decrease of 0.89 weeks worked (at the 1% level). These mixed findings suggest that anti-discrimination laws may lead employed women to work longer hours per week but in fewer weeks overall, reflecting a complex adjustment on the intensive margin.
@@ 4.2 Wages
As shown in Table 4, anti-discrimination laws significantly affect labor force participation and employment outcomes, thereby potentially altering the composition of the employed sample and raising endogeneity concerns for wage analyses. Table 5 shows regression results on the wage gap, using real hourly wages and annual earnings in 2019 dollars, measured by percentile ranks and computed jointly for men and women. As Chen and Roth (2024) argue, this robust, scale-invariant metric reduces the impact of outliers and reflects an individual’s relative wage position, allowing us to evaluate how anti-discrimination laws shift wage standings for individuals in same-sex partnerships.
@@
For men in same-sex partnerships, the increased employment rates reported in Table 4 are accompanied by significant wage gains. Specifically, anti-discrimination laws boost hourly wages by roughly 2.7 percentiles and annual earnings by 3 percentiles on average. Furthermore, these laws raise the probability of ranking above the 25th and 50th percentiles in the hourly wage distribution by 3.6 and 3.4 percentage points, respectively, although the effect at the 75th percentile is not significant. In the annual earnings distribution, anti-discrimination laws improve the likelihood of being in the upper tiers by 3.6, 5.2, and 3.7 percentage points at the 25th, 50th, and 75th percentiles, respectively. For women in same-sex partnerships, the wage distribution remains largely unaffected by anti-discrimination laws, as evidenced by statistically insignificant estimates across the board. This finding is consistent with the muted effects on employment outcomes observed in Table 4 for women.
These findings suggest that the wage gains for men in same-sex partnerships operate primarily through increased employment. New entrants to the labor force, individuals who were previously at the bottom of the wage distribution, enter at various points, thereby shifting the overall distribution upward.
@@ 4.3 Impact Dynamics
We estimate Equation (1) using event study regressions following Borusyak et al. (2021) to trace the evolution of the impact estimates. Figure 3 displays the results for extensive labor supply outcomes: Panel (a) shows labor force participation rates, while Panel (b) presents employment levels. For men in male same‐sex partnerships, both panels reveal an immediate increase in labor force participation and employment following the enactment of ENDAs, although this positive effect wanes over time. In contrast, we see no impacts on women in same-sex partnerships. The pre-treatment trends are largely parallel, supporting the plausibility of the parallel trends assumption, yet the post-treatment dynamics diverge sharply by gender. The evidence here suggests that ENDAs lead to a temporary boost in labor market entry for men in same-sex partnerships, while not affecting employment and participation rates for women in same-sex partnerships.
Figure 4 presents the event study regressions for male and female pay disparities. The results mirror the trends observed in Figure 3, with a significant wage jump occurring one year after the law’s enactment for men in same-sex partnerships. However, the difference between the one-year and four-year post‑enactment estimates is relatively modest, suggesting that the initial wage improvement for men stabilizes and remains persistent over time. We find no statistically significant effects on wage disparities for women in any post‑ENDA period. These findings indicate that while anti‑discrimination laws yield immediate and lasting wage benefits for men in same‑sex partnerships, the impact on wage outcomes for women in same-sex partnerships is insignificant.
@@ 5 Robustness Checks
To ensure the robustness of our findings on the labor market impacts of anti-discrimination laws, we conduct a series of sensitivity analyses addressing potential sources of bias. Specifically, we examine (i) the endogeneity of law adoption using state-level support for same-sex marriage as a proxy for public sentiment, (ii) the potential confounding influence of same-sex marriage legalization, (iii) the role of migration patterns, (iv) geographic heterogeneity between metropolitan and non-metropolitan areas, (v) reporting changes associated with ACS redesign, (vi) the impact of applying sample weights, and (vii) the validity of our identification strategy using placebo tests and alternative estimators.
@@ 5.2 Marriage Laws
Table 6 presents results restricting the sample to periods before same-sex marriage legalization. Although this approach removes nearly half of the same-sex partnership observations and reduces statistical power, the qualitative patterns remain robust. Specifically, the point estimates for both labor force participation and employment remain comparable, exceeding 3 percentage points, with labor force participation retaining statistical significance. However, the employment coefficient for men declines from 0.037 to 0.0319, becoming borderline insignificant (p-value = 0.102), and similar shrinkages are observed on the intensive margin. In contrast, for women in same-sex partnerships, the intensive margin estimates are slightly larger. Overall, this sensitivity analysis indicates that our identification strategy is not driven by concurrent same-sex marriage policies, as the observed effects of anti-discrimination laws on labor market outcomes for same-sex couples persist despite the reduced sample size.
@@ 5.3 Migration
Next, we account for the possibility that migration patterns could influence our results. If LGB individuals were disproportionately moving to areas with recently enacted anti‑discrimination laws, it could bias the treatment effects. To address this concern, we restrict our sample to individuals who had not moved in the past year and re‑estimate the models presented in Table 4. By focusing on a more stable population, we can better isolate the effects of anti‑discrimination laws on labor outcomes. The corresponding estimates, reported in Table 7, remain consistent with the full sample results, providing further robustness and suggesting that migration does not significantly alter our conclusions. Interestingly, the estimates for men in same‑sex partnerships are slightly larger for this subset, which may indicate that the positive employment effects are even more pronounced among those who are less mobile.

### v3

@@ 4.1 Extensive & Intensive Margins
In Table 4, we present regression results that examine the extensive and intensive margins of labor supply for same-sex couples. The coefficient $\beta_{1g}$ captures the effect of anti-discrimination laws on the labor supply outcomes for men in male same-sex couples and for women in female same-sex couples.
Table 4: Effect of Anti-Discrimination Laws: Extensive & Intensive Margin of Labor Supply
(1) (2) (3) (4) (5) (6)
Extensive Margin Intensive Margin
ENDA $\times$ SSP: Female -0.0037 -0.0050 0.9519 -0.4143 0.7017∗∗ -0.8915∗∗∗
(0.0116) (0.0135) (0.5332) (0.6235) (0.3080) (0.3349)
ENDA $\times$ SSP: Male 0.0379∗∗∗ 0.0370∗∗∗ 1.1289∗∗ 1.5089∗∗ -0.0920 -0.0719
(0.0130) (0.0135) (0.4984) (0.6276) (0.3727) (0.2821)
$p$-value: M - F 0.0032 0.0177 0.7908 0.0082 0.1160 0.0465
Employed Only X X
Observations 7,781,358 7,781,358 7,781,358 7,781,358 5,765,580 5,765,580
Mean (F) 0.818 0.782 34.42 39.51 41.25 47.27
Mean (M) 0.825 0.790 35.53 39.96 42.41 47.69
  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019. We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment. The regressions are estimated jointly, allowing the treatment effect to differ for men and women through interaction terms. Columns (1–2) focus on the extensive margin, examining changes in labor force participation and employment. Columns (3–6) explore the intensive margin using variation in weeks worked per year and usual weekly hours; Columns (5–6) restrict the sample to employed individuals. The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively. Reported $p$-values test whether the ENDA effect differs significantly between men and women. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
@@
The joint estimation allows the effects of anti-discrimination laws to differ for men and women through interaction terms. Columns (1) and (2) show that, for men in male same-sex couples, anti-discrimination laws reduce both the labor force participation and employment gaps by approximately 3.8 and 3.7 percentage points, which translates into reductions of about 78% and 77%, respectively. These coefficients are statistically significant at the 1% level. Columns (3) through (6) address the intensive margin, indicating modest increases in weekly hours and weeks worked among all men. However, when the sample is restricted to employed men (Columns (5) and (6)), these effects become statistically insignificant. This pattern suggests that the primary channel through which these laws operate is by increasing the rate of employment rather than by altering hours worked among those already employed.
The estimates for women in same-sex couples follow a different pattern. Along the extensive margin (Columns (1) and (2)), the coefficients are small, negative, and statistically insignificant, on the order of 0.3–0.5 percentage points, which are not only imprecise but also economically small relative to mean participation rates. On the intensive margin, Column (3) shows a modest increase of 0.95 weekly hours (significant at the 10% level), while Column (4) shows a small, statistically insignificant reduction in weeks worked. When focusing on employed women (Columns (5) and (6)), we observe a statistically significant increase of 0.70 weekly hours and a significant decrease of 0.89 weeks worked. These mixed but generally modest effects underscore that, for women, the evidence for labor-market responses is weaker and less systematic than for men, consistent with the possibility that adjustments for women occur outside the labor-market margins examined here.
Finally, the joint estimation allows for explicit comparison across sexes. The $p$-values reported in the table indicate that the extensive margin effects differ significantly between men and women, with men experiencing large, positive, and statistically significant gains while women show no measurable changes. Additionally, the intensive margin effect on weeks worked is also significantly larger for men than for women. These differences underscore meaningful heterogeneity in how anti-discrimination laws affect male and female same-sex couples in the labor market.
@@ 4.2 Wages
As shown in Table 4, anti-discrimination laws significantly affect labor force participation and employment outcomes, thereby potentially altering the composition of the employed sample and raising endogeneity concerns for wage analyses. Table 5 shows regression results on the wage gap, using real hourly wages and annual earnings in 2019 dollars, measured by percentile ranks and computed jointly for men and women with heterogeneous ENDA effects by sex. [^14] We first compute weighted within-year percentile ranks for real annual earnings and real hourly wages by assigning individuals to weighted percentiles each survey year; we then construct three distributional indicators, set to one when an individual’s percentile rank meets or exceeds the 25th, 50th, and 75th percentiles, respectively. This robust, scale-invariant metric reduces the impact of outliers and reflects an individual’s relative wage position, allowing us to evaluate how anti-discrimination laws shift wage standings for individuals in same-sex couples. [^15] Following chen2024logs, including non-employed individuals in the wage-rank distribution improves transparency by combining extensive- and intensive-margin responses into a single metric. However, this approach does not recover the causal effect of ENDAs on potential wages for all individuals, since ranks for non-employed individuals are mechanically assigned rather than observed wage offers.
Table 5: Effect of Anti-Discrimination Laws: Wages and Earnings
Hourly Wage Annual Earnings
(1) (2) (3) (4) (5) (6) (7) (8)
Percentile $\geq$25th $\geq$50th $\geq$75th Percentile $\geq$25th $\geq$50th $\geq$75th
ENDA $\times$ SSP: Female -0.3333 0.0045 -0.0117 -0.0061 -0.5994 0.0051 -0.0204 -0.0016
(0.8450) (0.0118) (0.0143) (0.0154) (0.8315) (0.0118) (0.0134) (0.0161)
ENDA $\times$ SSP: Male 2.6310∗∗ 0.0366∗∗∗ 0.0326 0.0232 2.9315∗∗∗ 0.0365∗∗∗ 0.0467∗∗ 0.0334∗∗
(1.1771) (0.0139) (0.0182) (0.0178) (1.1019) (0.0139) (0.0185) (0.0154)
$p$-value: M - F 0.0379 0.0340 0.0776 0.2093 0.0066 0.0382 0.0049 0.0713
Observations 7,781,358 7,781,358 7,781,358 7,781,358 7,781,358 7,781,358 7,781,358 7,781,358
Mean (F) 48.73 0.799 0.513 0.247 48.68 0.798 0.515 0.241
Mean (M) 51.51 0.791 0.557 0.305 51.69 0.791 0.561 0.305
  • Notes: Data are derived from the American Community Survey (ACS) conducted annually between 2005 and 2019. We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment. The regressions are estimated jointly, allowing the treatment effect to differ for men and women through interaction terms. Columns (1–4) report effects on hourly real wage ranks, while Columns (5–8) report effects on annual real earnings ranks. Columns (2–4) and (6–8) estimate the probability of being above the 25th, 50th, and 75th percentiles of the respective income distributions. All earnings measures are expressed in 2019 dollars. The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively. Reported $p$-values test whether the ENDA effect differs significantly between men and women. Standard errors are clustered at the county level.
  • $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
@@
For men in same-sex couples, the increased employment rates reported in Table 4 are accompanied by significant wage gains. Specifically, anti-discrimination laws boost hourly wages by roughly 2.6 percentiles and annual earnings by 2.9 percentiles on average. Furthermore, these laws raise the probability of ranking above the 25th and 50th percentiles in the hourly wage distribution by 3.7 and 3.3 percentage points, respectively, although the effect at the 75th percentile remains statistically insignificant. In the annual earnings distribution, anti-discrimination laws improve the likelihood of being in the upper tiers by 3.6, 4.7, and 3.3 percentage points at the 25th, 50th, and 75th percentiles, respectively. For women in same-sex couples, the wage distribution remains largely unaffected by anti-discrimination laws. The estimates are statistically insignificant and also small in magnitude—typically well under one percentile point, suggesting limited scope for economically meaningful wage effects. This pattern is consistent with the muted employment responses for women reported in Table 4.
Joint estimation allows us to formally compare effects across sexes. The $p$-values reported in Table 5 show that the wage effects for men are significantly larger than those for women at most points in the wage and earnings distributions, reinforcing that anti-discrimination laws generate substantial improvements in labor market standing for men in same-sex couples but have comparatively limited effects for women.
These findings suggest that the wage gains for men in same-sex couples operate primarily through increased employment on the extensive margin. New entrants to the labor force, individuals who were previously non-employed or concentrated at the bottom of the distribution, enter at various points in the wage distribution, shifting the overall distribution upward, particularly around the lower and middle percentiles where such workers are most likely to sort. Specialization models among couples imply that primary earners are typically already located high in the earnings distribution; thus, if ENDAs primarily induce secondary earners to enter work or increase their labor supply, we should expect relatively muted effects at the upper end of the hourly wage distribution. This is precisely what we observe: the 75th percentile of the hourly wage distribution shows no significant response to ENDAs. By contrast, the 75th percentile of the annual earnings distribution rises significantly. While this may appear surprising at first glance, it likely reflects a combination of wage gains, increased hours or weeks worked, and the fact that some new entrants can move into higher annual earnings ranks through expanded labor supply. Overall, the evidence points to selection through secondary earners entering employment and adjusting labor supply, rather than upward movements among already high-earning primary earners.
@@ 4.3 Impact Dynamics
We estimate Equation (1) using event study regressions following borusyak2021revisiting to trace the evolution of the impact estimates. Figure 3 displays the results for extensive labor supply outcomes: Panel (a) shows labor force participation rates, while Panel (b) presents employment levels. For men in male same-sex couples, both panels show an immediate rise in labor force participation and employment following ENDA adoption, with the largest effects appearing in the early post-treatment years. While the point estimates decline somewhat over time, our data do not extend far enough to determine whether these effects are temporary or simply leveling off. In contrast, the estimates for women in same-sex couples are small and statistically insignificant, indicating little evidence of a labor-market response. The pre-treatment trends are largely parallel, supporting the plausibility of the identifying assumption, and the post-treatment dynamics again highlight pronounced gender differences in responses to ENDAs.
((a)) Labor Force Participation
((b)) Employment
Figure 3: Labor Supply Impacts on the Extensive Margin
Notes: Event study plots on the difference in labor supply between people in same-sex couples and different-sex couples broken down by sex following the county-level and state-level anti-discrimination laws. Coefficients with 95% confidence intervals. Standard errors are clustered at the county level. Reference year is -4.
((a)) Percentile Rank of Real Hourly Wage
((b)) Percentile Rank of Annual Wage
Figure 4: Impacts on Wages
Notes: Event study plot on the difference in pay between people in same-sex couples and different-sex couples broken down by sex following the county-level and state-level anti-discrimination laws. Coefficients with 95% confidence intervals. Standard errors are clustered at the county level. Reference year is -4.
Figure 4 presents the event study regressions for male and female pay disparities. The results mirror the trends observed in Figure 3, with a significant wage jump occurring one year after the law’s enactment for men in same-sex couples. However, the difference between the one-year and four-year post‑enactment estimates is relatively modest, suggesting that the initial wage improvement for men stabilizes and remains persistent over time. We find no statistically significant effects on wage disparities for women in any post‑ENDA period. These findings indicate that while anti‑discrimination laws yield immediate and lasting wage benefits for men in same‑sex couples, the impact on wage outcomes for women in same-sex couples is insignificant.
@@ 5 Robustness Checks
To assess the robustness of our findings and address key threats to identification, we conduct several sensitivity analyses. These tests examine (i) the endogeneity of law adoption using state-level support for same-sex marriage, (ii) potential confounding from same-sex marriage legalization, (iii) migration responses, (iv) metro–non-metro heterogeneity, (v) sensitivity to commuting, and (vi) reporting changes from the ACS redesign.
For exercises (ii)–(vi), we estimate a fully interacted framework that splits men and women into subsample and complementary groups, allowing us to recover ENDA effects for each subgroup and test differences across sex and across samples.
@@ 5.1 Endogenous Adoption of Anti-Discrimination Laws
Figure 5: Impacts on Same-Sex Marriage Support
Notes: Event study plot showing how support for same-sex marriage changes following state-level anti-discrimination laws. Polling data comes from Pew Polling and encompasses 2005-2017 for every state except Hawaii and Alaska, which are missing for 2005-2008. Standard errors are clustered at the state level.
@@ 5.2 Marriage Laws
Table 6: Effects of Anti-Discrimination Laws on Labor Supply: Pre–Same-Sex Marriage Legalization
(1) (2) (3) (4) (5) (6)
Extensive Margin Intensive Margin
Pre–SSM Legalization
ENDA $\times$ SSP: Female -0.0157 -0.0215 1.1714 -0.6654 1.0376∗∗∗ -1.2330∗∗∗
(0.0153) (0.0158) (0.6726) (0.7928) (0.3682) (0.3997)
ENDA $\times$ SSP: Male 0.0283 0.0311 0.8892 0.4945 -0.1655 -0.7909∗∗
(0.0186) (0.0193) (0.6512) (0.8727) (0.4310) (0.3950)
$p$: M – F (Pre–SSM) 0.0186 0.0202 0.7304 0.2214 0.0192 0.4193
$p$: M (Pre–SSM – Post-SSM) 0.1196 0.7317 0.0293 0.0048 0.0000 0.0004
$p$: F (Pre–SSM – Post-SSM) 0.0702 0.0039 0.0024 0.0270 0.0299 0.0125
Observations 7,779,108 7,779,108 7,779,108 7,779,108 5,763,735 5,763,735
Mean (F) 0.817 0.774 34.44 39.13 41.28 46.85
Mean (M) 0.818 0.778 35.39 39.44 42.40 47.27
  • Notes: Data are derived from the American Community Survey (ACS) conducted annually between 2005 and 2019. We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment. All regressions in this table are estimated jointly with a fully interacted model that allows ENDA effects to vary across four groups: men and women, each in the pre–same-sex-marriage and post–same-sex-marriage periods. This framework nests the separate-by-sex estimations and enables formal tests of equality across groups. We report the ENDA effects estimated on the Pre–SSM Legalization sample and include $p$-values testing whether these coefficients differ from those obtained in the complementary Post–SSM sample. Columns (1–2) focus on the extensive margin, examining changes in labor force participation and employment. Columns (3–6) analyze intensive-margin outcomes, usual weekly hours and weeks worked, where Columns (5–6) restrict the sample to employed individuals. The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples in the Pre-SSM sample, respectively. Reported $p$-values test differences across sex and across robustness conditions. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 6 presents results from the jointly estimated model, reporting ENDA effects separately for the pre–same-sex-marriage subsample while also providing $p$-values testing whether these effects differ significantly from those in the complementary post-legalization period. Although this approach removes nearly half of the same-sex partnership observations and reduces statistical power, the qualitative patterns remain robust. For men in same-sex couples, the point estimates for both labor force participation and employment remain comparable to the baseline, exceeding 3 percentage points, with labor force participation retaining statistical significance. The employment coefficient declines from 0.037 to 0.0311 and becomes borderline insignificant, consistent with the loss of power, and similar attenuations appear on the intensive margin. For women, the intensive-margin estimates are slightly larger in the restricted sample, though still small in magnitude and statistically weak.
The pre–post $p$-value tests reported in Table 6 indicate that several coefficients differ across the two periods, particularly on the intensive margin. Still, the qualitative pattern of the ENDA effects is preserved: the signs of the coefficients remain stable, and the largest impacts continue to appear on the extensive margin for men and on the intensive margin for women. Moreover, even within the pre–SSM subsample, the male–female comparison on the extensive margin remains statistically significant and positive, reinforcing that men in same-sex couples experience systematically larger gains in labor force participation and employment. Thus, while same-sex marriage legalization may shift the magnitudes of some estimates, it does not overturn the core conclusion that anti-discrimination laws improve labor market outcomes for individuals in same-sex couples. This confirms that our identification strategy is not driven solely by concurrent same-sex marriage policies.
@@ 5.3 Migration
Next, we account for the possibility that migration patterns could influence our results. If LGB individuals were disproportionately moving to areas with recently enacted anti-discrimination laws, it could bias the treatment effects. To address this concern, we restrict our sample to individuals who had not moved in the past year and re-estimate the models presented in Table 4. By focusing on a more stable population, we can better isolate the effects of anti-discrimination laws on labor outcomes. The corresponding estimates, reported in Table 7, remain consistent with the full sample results, providing further robustness and suggesting that migration does not significantly alter our conclusions. Interestingly, the estimates for men in same-sex couples are slightly larger for this subset, which may indicate that the positive employment effects are even more pronounced among those who are less mobile.
Table 7: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of Non-Movers
(1) (2) (3) (4) (5) (6)
Extensive Margin Intensive Margin
Non-Movers
ENDA $\times$ SSP: Female 0.0050 0.0048 1.1290 0.1140 0.7276∗∗ -0.5031
(0.0142) (0.0147) (0.6189) (0.7471) (0.3084) (0.3944)
ENDA $\times$ SSP: Male 0.0421∗∗∗ 0.0424∗∗∗ 1.2112∗∗ 1.7705∗∗ -0.0583 0.2447
(0.0145) (0.0153) (0.5690) (0.7291) (0.4024) (0.2941)
$p$: M – F (Non-Movers) 0.0209 0.0412 0.9143 0.0503 0.1284 0.1084
$p$: M (Non-Movers – Movers) 0.3954 0.2811 0.6663 0.5295 0.9676 0.4559
$p$: F (Non-Movers – Movers) 0.0504 0.0763 0.2704 0.0013 0.2691 0.0146
Observations 7,781,169 7,781,169 7,781,169 7,781,169 5,765,391 5,765,391
Mean (F) 0.815 0.783 34.34 39.57 41.32 47.55
Mean (M) 0.823 0.791 35.44 39.99 42.45 47.89
  • Notes: Data are derived from the American Community Survey (ACS) conducted annually between 2005 and 2019. We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment. All regressions in this table are estimated jointly with a fully interacted model that allows ENDA effects to vary across four groups: men and women, each in the mover subsample and in the non-mover subsample. This framework nests the separate-by-sex estimations and enables formal tests of equality across groups. We report the ENDA effects estimated for non-movers and include $p$-values testing whether these coefficients differ from those obtained in the complementary sample of recent movers. Columns (1–2) focus on the extensive margin, examining changes in labor force participation and employment. Columns (3–6) analyze intensive-margin outcomes, usual weekly hours and weeks worked, where Columns (5–6) restrict the sample to employed individuals. The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples in the non-mover sample, respectively. Reported $p$-values test differences across sex and across the mover versus non-mover samples. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
The joint-estimation framework, which interacts ENDA exposure with both gender and migration status (male–mover, male–non-mover, female–mover, female–non-mover), also allows us to test whether the ENDA effects differ between movers and non-movers. For men, the $p$-value tests show no statistically significant differences across the two samples, reinforcing that migration does not meaningfully drive the male results. Among women, several coefficients do differ between movers and non-movers, but the overall qualitative pattern of effects is preserved and most estimates remain small or statistically insignificant. Importantly, the male–female comparison on the extensive margin remains statistically significant within the non-mover subsample, indicating that the gender differences highlighted in the main specification are not driven by migration behavior.